OpenStudy (anonymous):

Will Fan And Medal <3 7. Simplify the radical expression. 2 square root 6 + 3 square root 96 A) 14 square root 6 B) 14 square root 96 C) 5 square root 96 D) 50 square root 6 8. Simplify the radical expression. (8 + square root 11)(8 – square root 11) A) 53 B) 75 + 16 sqrt 11 C) -57 D) 64 + sqrt 11 9. What is the domain of the function? Y = sqrt 4x+8 A) x > - 2 (> is underlined) B) x < - 2 (< is underlined) C) x > 2 (> is underlined) D) x < 2 (< is underlined)

OpenStudy (elvisiscool69):

for a test?

OpenStudy (anonymous):

does it help to know that \(96=16\times 6\)?

OpenStudy (anonymous):

@elvisiscool69 Practice questions.

OpenStudy (anonymous):

the reason that is useful is because that means \[\sqrt{96}=\sqrt{16\times 6}=\sqrt{16}\times \sqrt6=4\sqrt6\]

OpenStudy (anonymous):

@satellite73 Are you referring to question #7

OpenStudy (anonymous):

making the first question \[2\sqrt6+3\times 4\sqrt6\]

OpenStudy (anonymous):

yes, the one with the 96 in it

OpenStudy (anonymous):

@satellite73 I got B

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

what is \(3\times 4\)?

OpenStudy (anonymous):

lol don't let anyone bust your chops about a profile picture i am not really a bike either you can be anyone you like

OpenStudy (anonymous):

12

OpenStudy (anonymous):

ok so now we have \[2\sqrt6+12\sqrt6\] and we have to add

OpenStudy (anonymous):

@MariahDimond no, i am a car

OpenStudy (anonymous):

in fact, it am a 73 satellite (google it)

OpenStudy (anonymous):

@satellite73 I'm not sure what else to do..

OpenStudy (anonymous):

@Answers101 so what is 2 apples plus twelve apples?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

14 whats?

OpenStudy (anonymous):

sure did you google it? @Answers101 2 cars plus 12 cars is 14 whats?

OpenStudy (anonymous):

So the answer is A

OpenStudy (anonymous):

oui

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@satellite73 OK question 8 I got D... is that correct?

OpenStudy (anonymous):

no, D is wrong

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\\ (8+\sqrt{11})(8-\sqrt{11})=8^2-11\]

OpenStudy (anonymous):

@MariahDimond a picture of your picture

OpenStudy (anonymous):

@satellite73 yay you found me :D

OpenStudy (anonymous):

@satellite73 Is it A?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@satellite73 9. What is the domain of the function? Y = sqrt 4x+8 A) x > - 2 (> is underlined) B) x < - 2 (< is underlined) C) x > 2 (> is underlined) D) x < 2 (< is underlined)

OpenStudy (anonymous):

you cannot take the square root of a negative number set \[4x+8\geq 0\] and solve for \(x\) in two steps

OpenStudy (anonymous):

@satellite73 I think the answer is C

OpenStudy (anonymous):

nope